OpenStudy (anonymous):

Consider the point (x,y) lying on the graph of the line 2x+4y=5. Let L be the distance from point (x,y) to the origin. Write L as a function of x

OpenStudy (anonymous):

Well, we know: \[ d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2} \]And we know the origin is at (0,0). So, this simplifies to: \[ d=\sqrt{x_1^2+y_1^2} \]Try to write y as a function of x, and substitute it into the formula.

OpenStudy (anonymous):

I got d=sqrt(x^2+((5-2x)^2/4^2))

OpenStudy (anonymous):

Yeah, now expand and simplify, and et voilá

OpenStudy (anonymous):

Stuck at \[d =\sqrt{x ^{2}+\left( \left( 5-2x \right)^{2}\div16 \right)}\]

OpenStudy (anonymous):

It's an ugly expression, I don't know if it is a requirement to simplify, but the answer you gave is correct.

OpenStudy (anonymous):

The book says its \[L = 20x ^{2}+20x +25\]

OpenStudy (anonymous):

I mean, we have, simplifications and all \[ L=\frac{\sqrt{20x^2-20x+25}}{4} \]But I don't see how the book ended without a fraction.

OpenStudy (anonymous):

Thats what I said! But one last thing, I don't understand how the 4 got as the denominator

OpenStudy (anonymous):

Since: \[ \sqrt{16}=4 \]We can simply remove it from the radical and use the latter value. Sorry for responding so late! For some reason or another I can't see this thread's replies!